OpenStudy (anonymous):

Factor x2 - 8x + 9. (4 points) (x - 9)(x - 1) (x + 9)(x + 1) Prime (x + 9)(x - 1)

OpenStudy (compassionate):

\[x^2 - 8x + 9\] So, basically, what factors when added have a sum of 8 and a product of 9?

OpenStudy (compassionate):

\[(x \pm ?)(x \pm ?)\]

OpenStudy (anonymous):

x+9 x- 1 ??

OpenStudy (compassionate):

\[(x + 9)(x - 1) = x^2 - x + 9x - 9 = x^2 + 8x - 9\]

OpenStudy (anonymous):

._.

OpenStudy (compassionate):

8 and 9 have no summation and products. There, it is prime.

OpenStudy (anonymous):

oh.... Thanks!